You have been given the following information about the production of Usher Co., and are asked to provide the plant manager with information for a meeting with the vice president of operations. Standard Cost Card Direct materials (5 pounds at $5 per pound) $25.00 Direct labor (0.90 hours at $10) 9.00 Variable overhead (0.90 hours at $4 per hour) 3.60 Fixed overhead (0.90 hours at $9 per hour) 8.10 $45.70 The following is a variance report for the most recent period of operations. Variances Costs Total Standard Cost Price Quantity Direct materials $405,000 $8,298 F $9,900 U Direct labor 145,800 4,590 U 7,200 U

A.) How many units were produced during the period?

B.) How many pounds of raw materials were purchased and used during the period?

C.) What was the actual cost per pound of raw materials?

D.) How many actual direct labor hours were worked during the period?

E.) What was the actual rate paid per direct labor hour?
Answer #1

a) Unit produced = 405000/25 = 16200 Units

b) Material quantity variance = (Standard quantity-actual quantity)Standard price

-9900 = (16200*5*5-5X)

5X = 414900

X(actual pound) = 414900/5 = 82890 Pounds

c) Material price variance = (Standard price-actual price)actual quantity

8298 = (5*82980-82980X)

82980X = 406602

X(actual Cost per pound) = 4.9 per pound

d) Labor efficiency variance = (Standard hour-actual hour)Standard rate

-7200 = (16200*.9-X)*10

-7200 = 145800-10X

10X = 153000

X(actual hour) = 15300 Hour

e) Labor rate variance = (Standard rate-actual rate)actual hour

-4590 = (10*15300-15300X)

15300X = 153000+4590

X(actual rate) = 10.3 per hour